# Efficient LiDAR Reflectance Compression via Scanning Serialization

### v1

@@ Appendix E Downstream Tasks
Appendix F More Ablation Studies
Sequence Parallelization. In Section 3.4 – Dual Parallelization, we propose the dual parallelization scheme. including sequence level and window level, to accelerate coding speed of SerLiC within a frame. The ablation of window-level parallelization is presented in Section 4.5 – Window-based Parallelization. Limited by page number, we depict the ablation study of Sequence-level Parallelization here.
@@ Appendix F More Ablation Studies
Fig. LABEL:fig:supp_seq_corre visualizes the reflectance of a point cloud frame. For better observation, we color the reflectance values using different colors. It is clear that reflectance values are close to each other in a scan line while different to some extent across scan lines. This intuitively reveals that using point correlations within a scan line leads to better performance. Furthermore, we conduct experiments which exploit correlations not only across points in a sequence but also across sequences for reflectance compression. In this way, the sequence-level parallelization is disabled. That is, we apply the autoregressive coding both across sequences (called inter-sequence coding in the following) and within a sequence (called intra-sequence coding in the following).
@@
Computational Complexity. As the inter-sequence coding disable sequence-level parallelism, its memory cost is reduced. As reported in Table 7, the memory usage of inter-sequence coding us only 0.04 GB. However, the runtime is accordingly remarkably increased due to the autoregressive coding across sequences. For example, the total encoding/decoding time is 5.20/9.84 seconds per frame when using the previous sequence as context of the current sequence. When using three sequences, the coding time is even longer (5.48/10.16) as more times are required by the context construction stage.

### v2

@@ Appendix D Downstream Tasks
Appendix E More Ablation Studies
Sequence Parallelization. In Section 3.4—Dual Parallelization, we propose the dual parallelization scheme. including sequence level and window level, to accelerate coding speed of SerLiC within a frame. The ablation of window-level parallelization is presented in Section 4.5—Window-based Parallelization. Limited by page number, we depict the ablation study of Sequence-level Parallelization here.
@@ Appendix E More Ablation Studies
LABEL:fig:supp_seq_corre visualizes the reflectance of a point cloud frame. For better observation, we color the reflectance values using different colors. It is clear that reflectance values are close to each other in a scan line while different to some extent across scan lines. This intuitively reveals that using point correlations within a scan line leads to better performance. Furthermore, we conduct experiments which exploit correlations not only across points in a sequence but also across sequences for reflectance compression. In this way, the sequence-level parallelization is disabled. That is, we apply the autoregressive coding both across sequences (called inter-sequence coding in the following) and within a sequence (called intra-sequence coding in the following).
@@
Computational Complexity. As the inter-sequence coding disable sequence-level parallelism, its memory cost is reduced. As reported in Table 7, the memory usage of inter-sequence coding is only 0.04 GB. However, the runtime is accordingly remarkably increased due to the autoregressive coding across sequences. For example, the total encoding/decoding time is 5.20/9.84 seconds per frame when using the previous sequence as context of the current sequence. When using three sequences, the coding time is even longer (5.48/10.16) as longer time is required by the context construction stage.